OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope? Point: (–3, 7); Slope: 4 A. y – 7 = 4(x – 3) B. y – 7 = 4(x + 3) C. y + 7 = 4(x + 3) D. y + 7 = 4x – 12

OpenStudy (anonymous):

@tgawade

OpenStudy (anonymous):

Same as previous use slope point form y - y1 = m * (x - x1)

OpenStudy (anonymous):

y – 7 = 4(x + 3)

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Correct !!